OpenStudy (anonymous):

What is the range of the following function and how do u get it? 1/sqrt 1+x

OpenStudy (anonymous):

to ge tthe range u must get the graph or find the inverse

OpenStudy (anonymous):

is it \[f(x)=\frac{1}{\sqrt{x+1}}\] ?

OpenStudy (anonymous):

it should be clear that \(\sqrt{x+1}\geq0\) so this can never be negative, in other words \[\frac{1}{\sqrt{x+1}}>0\] for all \(x\) it can never be zero, because a fraction is only zero if there is a zero in the numerator so the range is \((0,\infty)\)

OpenStudy (anonymous):

i guess one more thing needs to be said: the closer \(x\) gets to \(-1\) the smaller \(x+1\) gets, and therefore the larger \(\frac{1}{\sqrt{x+1}}\) gets that is why it goes to infinity

OpenStudy (anonymous):

thamks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

How would you do the following initial value problem d^2y/dt^2 + dy/dt = 2t where t is greater then 0 and y(0) = 0 and dy/dx (0) = 1